OpenStudy (anonymous):

a roll of wallpaper covers an area of 28ft square. If the roll is 1 17/24 wide, how long is the roll?

OpenStudy (anonymous):

\[28 \times\frac{24}{41} = 16.39ft\]

OpenStudy (anonymous):

im confuse where did you get the 24/41? can you show me?

OpenStudy (anonymous):

\[1 \frac{17}{24}ft = \frac{41}{24}ft\] \[width \times height = 28ft^{2}\] Substitute width with what it gave you(41/24)\[\frac{41}{24}ft \times height = 28ft^{2}\]then divide both sides by (41/24). When you divide by a fraction, you take its reciprocal and multiply it:\[height = 28ft^{2} \times \frac{24}{41}ft^{-1}\]

OpenStudy (anonymous):

thank you it makes way more sense now

OpenStudy (anonymous):

No problem. A medal would be nice :) Just click the "Good Answer" button next to my name

OpenStudy (anonymous):

\[28 ft/roll \times 41/24 ft roll = 287/6 or 47.33\]